Business and finance. When x number of units are sold, the price of each unit (in dollars) is given by p=-x/2+75. Find the unit price when the following quantities are sold: 2, 7, 9, 11.
Answer by checkley75(3666)

P=-X/2+75
P=-2/2+75
P=-1+75
P=74 ANSWER
P=-7/2+75
P=-3.5+75
P=71.5 ANSWER.
P=-9/2+75
P=-4.5+75
P=70.5 ANSWER.
P=-11/2+75
P=-5.5+75
P=69.5 ANSWER.